OpenStudy (anonymous):

I need help understanding how to find the angle between vectors! u = <2, -4>, v = <3, -8> Someone to walk me through it would be great!

OpenStudy (zpupster):

\[\cos \alpha =\frac{ (u1v1 + u2v2) }{ \sqrt{u1^2 +{ u2^2}} {\sqrt{v1^2 +v2^2}} }\]

OpenStudy (anonymous):

So, ((2*3)+(-4*-8)) /sqrt of (2)^2+(-4)^2 * sqrt of (3)^2+(-8)^2

OpenStudy (anonymous):

I need to do the math really quick.. one sec..

OpenStudy (zpupster):

lets plug them in (2)(3) + (-4)(-8)/(sqrt((2)^2)+(3^2)) (sqrt((-4^2)+(-8^2)) (6+32)/ ()sqrt(4+9)(sqrt(16 + 64)

OpenStudy (anonymous):

19(sqrt of 65)/130